OpenStudy (anonymous):

I'm confused on how to solve this...Find the limit of the function by using direct substitution. limit as x approaches quantity pi divided by two of quantity three times e to the x times cosine of x.

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

Here's a hint...since everything is multiplied together.. what is \(\large \cos(\pi/2)\)?

OpenStudy (anonymous):

0

OpenStudy (johnweldon1993):

right...and since 0is being multiplied, we know the whole function approaches 0 as x goes to pi/2

OpenStudy (anonymous):

where do i plug in 0

OpenStudy (anonymous):

\[\large 3e^\frac{\pi}{2}\cos(\frac{\pi}{2})\] is what is meant by direct substitution i.e. plug it in

OpenStudy (anonymous):

so it's 3e^0*0 and that gives 0

OpenStudy (johnweldon1993):

No, the 'x' you plug in is \(\large \frac{\pi}{2}\) so really it would be as satellite put above \[\large 3e^{\frac{\pi}{2}}cos(\frac{\pi}{2})\] And since we know that \(\large cos(\frac{\pi}{2}) = 0\) we have \[\large 3e^{\frac{\pi}{2}} \times 0\] which we know will be 0 since anything times 0 = 0

OpenStudy (anonymous):

ohhh okay got it thanks :)

OpenStudy (johnweldon1993):

Anytime! :)